# Onsager reciprocal relations for charge and spin transport in periodically driven systems

### v1

@@ VI.1 Origin of the $\theta$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$
×$\times[e^{-ik_{y}}(-1)^{n-m-(\beta-1)l}+e^{ik_{y}}],$ (101)
and $\mathcal{J}_{l}(u)$ is the $l$th-order Bessel function of the first kind. For simplicity, we restrict the Floquet indices $m$, $n$, and $l$ to be $-1\leq m$, $n$, $l\leq 1$, which corresponds to the case at $n_{\textrm{max}}=1$. After some calculation, we obtain $I_{x}^{mn}(k_{x},u,\theta,\beta)$’s at $\beta=2$ and $3$,
$I_{x}^{mn}(k_{x},u,\theta,2)$
$\sim 2\delta_{n-m,0}\cos k_{x}\mathcal{J}_{0}(u)^{2}$
@@
$-2(e^{-i\theta}\delta_{n-m,2}+e^{i\theta}\delta_{n-m,-2})\sin k_{x}\mathcal{J}_{0}(u)\mathcal{J}_{1}(u)+O(u^{3}),$ (102)
@@
$I_{x}^{mn}(k_{x},u,\theta,3)$
$\sim 2\delta_{n-m,0}\cos k_{x}\mathcal{J}_{0}(u)^{2}$
@@
$-2(e^{-i\theta}\delta_{n-m,2}+e^{i\theta}\delta_{n-m,-2})\cos k_{x}\mathcal{J}_{1}(u)^{2}+O(u^{3}).$ (103)
@@
$I_{x}^{mn}(k_{x},u,\theta,2)$
$\sim\delta_{n-m,0}(2-u^{2})\cos k_{x}-(\delta_{n-m,1}+\delta_{n-m,-1})u\sin k_{x}$
@@
$-(e^{-i\theta}\delta_{n-m,2}+e^{i\theta}\delta_{n-m,-2})u\sin k_{x}+O(u^{3}),$ (104)
@@
$I_{x}^{mn}(k_{x},u,\theta,3)$
$\sim\delta_{n-m,0}(2-u^{2})\cos k_{x}-(\delta_{n-m,1}+\delta_{n-m,-1})u\sin k_{x}$
@@
$-(e^{-i\theta}\delta_{n-m,2}+e^{i\theta}\delta_{n-m,-2})\frac{u^{2}}{2}\cos k_{x}+O(u^{3}).$ (105)
@@
These equations can explain the $\theta$-independent $\sigma_{yx}^{\textrm{S}}$ for small $u$ because the terms for $n-m=0$ are primary (if they give the finite contribution). Although the terms for $n-m\neq 0$ are necessary for $\sigma_{yx}^{\textrm{C}}\neq 0$, these equations can also explain the $\theta$-independent $\sigma_{yx}^{\textrm{C}}$ for small $u$, at which the terms for $n-m=\pm 1$ are dominant. This is because in the small-$u$ region, the $\theta$-independent $u^{1}$ terms are more important than the $\theta$-dependent $u^{2}$ terms. For moderately large $u$, the $\theta$-dependent $u^{2}$ terms for $n-m=\pm 1$ may become non-negligible. Therefore, the $\theta$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ could be interpreted in terms of the $u$ and $\theta$ dependences of the light-induced corrections to the kinetic energy.
VI.2 Comparisons with other relevant studies
@@
We now compare our results about the Onsager reciprocal relations with other relevant studies. As described in Sec. I, there are several studies Onsager-PD1 ; Onsager-PD2 ; Onsager-PD3 ; NA-Onsager about the Onsager reciprocal relations in periodically driven systems. In two of them Onsager-PD1 ; Onsager-PD2 , the time-periodic field was treated as perturbation. Meanwhile, in our theory, the time-periodic pump field has been treated nonperturbatively. In general, the time-periodic field should be treated nonperturbatively to describe periodically driven systems because its nonperturbative effects often change the electronic states drastically. Therefore, the systems studied in these papers Onsager-PD1 ; Onsager-PD2 are insufficient to discuss the standard periodically driven systems. Then, in another Onsager-PD3 , the Onsager reciprocal relations about the matter or energy transport between a periodically driven system and one of the three reservoirs were studied. This is distinct from the transport phenomena within periodically driven systems. Therefore, except our previous paper NA-Onsager , there has been no study of the Onsager reciprocal relations for any transport within periodically driven systems in which the nonperturbative effects of the driving field are taken into account. In that paper NA-Onsager , we numerically studied the time-averaged charge off-diagonal dc conductivities $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{xy}^{\textrm{C}}$ in graphene driven by CPL or LPL and found that $\sigma_{yx}^{\textrm{C}}=-\sigma_{xy}^{\textrm{C}}$ or $\sigma_{yx}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}$ holds with CPL or LPL, respectively. However, we did not discuss the generality of these relations, i.e., we just showed that they are satisfied in our numerical results. Meanwhile, in this paper, we have made general arguments about the Onsager reciprocal relations for spin and charge transport within the periodically driven systems. These arguments are applicable to many periodically driven systems and extendable to others. Therefore, this work is the first systematic study of the Onsager reciprocal relations in periodically driven systems. Most importantly, this paper is the first work demonstrating the Onsager reciprocal relations for the spin transport in periodically driven systems.
@@
Next, we compare our results in the case of BCPL at $\beta=3$ with our previous study NA-BCPL for graphene driven by BCPL at $\beta=3$. (Although we have also studied the case at $\beta=2$ there NA-BCPL , its results are not directly comparable to the results obtained in this paper due to the vital role of the valley degree of freedom in that case studied in Ref. NA-BCPL, .) There are three similarities between our results shown in Sec. V B 2 and those obtained in this previous study NA-BCPL : $\sigma_{yx}^{\textrm{C}}$ is almost independent of $\theta$ for small $u$ and depends on it for moderately large $u$; $\sigma_{yx}^{\textrm{C}}$ for small $u$ is dominated by the antisymmetric part; and the antisymmetric and symmetric parts are mixed for moderately large $u$ at $\theta=\frac{\pi}{2}$. Therefore, these properties may be characteristic properties of $\sigma_{yx}^{\textrm{C}}$ in systems driven by BCPL. Then, the difference is that the antisymmetric part of $\sigma_{yx}^{\textrm{C}}$ becomes almost vanishing for moderately large $u$ in Sr2RuO4 driven by BCPL at $\theta=0$ and $\pi$, whereas the antisymmetric part remains dominant even for moderately large $u$ in graphene driven by BCPL at these $\theta$’s NA-BCPL .
@@
We turn to the comparisons with some theoretical studies BCPL-pert1 ; BCPL-pert2 ; BCPL-pert3 about transport properties with BCPL at $\beta=2$. In these studies, the effects of the BCPL on charge or spin transport have been analyzed in perturbation theories. First, our results about the effects of the SOC are similar to those on the charge and spin conductivities of another three-orbital electron system having the $z$ component of the SOC in the second-order perturbation theory against the BCPL field BCPL-pert2 . This similarity may be reasonable because the SOC is treated nonperturbatively in both cases. Then, our $\theta$-dependent $\sigma_{yx}^{\textrm{C}}$ is similar to the $\theta$-dependent charge conductivity obtained in the third-order perturbation theory against the BCPL field BCPL-pert3 . However, a critical value of $u$ above which $\sigma_{yx}^{\textrm{C}}$ depends on $\theta$ exists only in our case [see Fig. 12(a)]. This critical value may arise from the nonlinear $u$ terms of the light-induced corrections, as we have discussed in Sec. VI A. Therefore, the crossover between the $\theta$-independent and $\theta$-dependent $\sigma_{yx}^{\textrm{C}}$ may be characteristic of nonperturbative effects of BCPL. This interpretation remains valid in graphene driven by BCPL NA-BCPL . Although the authors of Ref. BCPL-pert3, have claimed that a nonperturbative effect is also discussed, their discussions are insufficient because they have used an approximation for the Hamiltonian, $H_{0}[k+A(t)]\approx H_{0}(k)+A(t)\frac{\partial H_{0}(k)}{\partial k}$ [see Eq. (29) of Supplemental Material of Ref. BCPL-pert3, ], which is valid only if $A(t)$, the vector potential of the BCPL, can be treated perturbatively. Namely, their theory can analyze only the perturbative effects of the BCPL. We do not use such an approximation; instead, we have analyzed the nonperturbative effects of BCPL in the standard Floquet linear-response theory Tsuji ; Mikami ; Eckstein ; NA-FloquetSHE .
VI.3 Experimental realization
@@
Finally, we comment on experimental realization of our results. We have supposed that our periodically driven open system can reach a nonequilibrium steady state due to the damping $\Gamma$. The AHE predicted in such a periodically driven open system Oka-PRB was experimentally observed Light-AHE-exp2 . Therefore, our time-averaged charge off-diagonal dc conductivities could be experimentally observed in the pump-probe measurements for periodically driven Sr2RuO4. In addition, our SHE could be detected via the inverse SHE InvSHE1 ; InvSHE2 because we have demonstrated that $\sigma_{yx}^{\textrm{S}}$ satisfies the Onsager reciprocal relation, which is similar to that in nondriven systems, and is dominated by the antisymmetric part. In the case of Sr2RuO4, $u=0.1$ for $\Omega=6$ eV corresponds to $E_{0}\approx 15.4$ MV/cm, where we have used $u=ea_{\textrm{NN}}A_{0}=ea_{\textrm{NN}}E_{0}/\Omega$ and $a_{\textrm{NN}}\approx 0.39$ nm Sr2RuO4-aNN . From an experimental point of view, the pump field of the order of $10$ MV/cm can be realized Iwai-review . Therefore, it may be possible to observe the $\theta$-dependent $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ for moderately large $u$ as well as the $\theta$-independent $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ for small $u$. Although there is a possibility that the similar $\theta$-dependent $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ will be realized even for smaller $\Omega$ and $E_{0}$, it is difficult to check this possibility due to the huge cost of the numerical calculations. Note that in the case of CPL NA-FloquetSHE , the $u$ dependence of the $\sigma_{yx}^{\textrm{S}}$ or $\sigma_{yx}^{\textrm{C}}$ obtained for $\Omega=6$ eV is similar to that obtained for smaller $\Omega$.
VII Conclusions
@@
Our arguments and numerical calculations have demonstrated that $\sigma_{yx}^{\textrm{S}}$ in the periodically driven systems satisfies the Onsager reciprocal relation. This means that the finite $\sigma_{yx}^{\textrm{S}}$ in the periodically driven systems can be indirectly observed by measuring the time-averaged dc conductivity of the inverse SHE. Therefore, our results open the way for detecting the spin current in periodically driven systems via the inverse SHE. Since such an indirect detection method has been widely used in many spintronics phenomena of nondriven systems Bauer , this achievement is useful to develop and observe many spintronics phenomena in periodically driven systems. Therefore, our results provide a vital step towards comprehensive understanding and further development of spintronics in periodically driven systems.
@@
Moreover, our arguments have resolved the contradictory statements about the Onsager reciprocal relation for the spin transport in nondriven systems IncorrectOnsager-JS1 ; IncorrectOnsager-JS2 ; CorrectOnsager-JS1 ; CorrectOnsager-JS2 . There is a previous study claiming the violation of the Onsager reciprocal relation for the spin transport IncorrectOnsager-JS2 , whereas there are other studies claiming its existence IncorrectOnsager-JS1 ; CorrectOnsager-JS1 ; CorrectOnsager-JS2 . Furthermore, among the latter studies, the expression of the Onsager reciprocal relation is different: in one of them IncorrectOnsager-JS1 the spin off-diagonal dc conductivity is dominated by the symmetric part, whereas in the others CorrectOnsager-JS1 ; CorrectOnsager-JS2 it is dominated by the antisymmetric part. As we have shown in Sec. II B 1, our results are consistent with the results of the last two studies CorrectOnsager-JS1 ; CorrectOnsager-JS2 , i.e., the spin off-diagonal dc conductivity satisfies the Onsager reciprocal relation even with time-reversal symmetry and is dominated by the antisymmetric part. Therefore, our arguments support the validity of the interpretation of the inverse SHE InvSHE1 ; InvSHE2 as the existence of the spin current in nondriven systems.
@@
Then, our numerical calculations in the cases with BCPL have indicated that $\sigma_{yx}^{\textrm{C}}$ cannot necessarily be regarded as the anomalous Hall conductivity even if time-reversal symmetry is broken. In general, we can regard $\sigma_{yx}^{\textrm{C}}$ as the anomalous Hall conductivity if and only if $\sigma_{yx}^{\textrm{C}}$ is dominated by the antisymmetric part. In addition, time-reversal symmetry can be broken by BCPL NA-BCPL . As shown in Sec. V B 2, $\sigma_{yx}^{\textrm{C}}$ with BCPL for moderately large $u$ is almost vanishing or dominated by the symmetric part, although that for small $u$ is dominated by the antisymmetric part. This unusual property does not contradict the Onsager reciprocal relation because this relation in the case with BCPL restricts $\sigma_{yx}^{\textrm{C}}$ to neither the antisymmetric nor symmetric part generally (see Sec. II A 2). It may be due to the lack of a simple relation between the pump field of BCPL and its time-reversal counterpart. In contrast, the pump field of CPL has the simple relation $\mathbf{A}_{\textrm{LCPL}}(-t)=\mathbf{A}_{\textrm{RCPL}}(t)$; as a result, $\sigma_{yx}^{\textrm{C}}$ with CPL is restricted to the antisymmetric part (see Sec. II A 2). These results suggest that even if time-reversal symmetry can be broken by the pump field, it is highly required to check whether or not the main term of $\sigma_{yx}^{\textrm{C}}$ is given by the antisymmetric part in discussing the AHE. It is also necessary to check the main term of $\sigma_{yx}^{\textrm{S}}$ to discuss the SHE because in some cases such as the cases with BCPL its main term is not restricted by the Onsager reciprocal relation to either the antisymmetric or symmetric part. These suggestions are useful for future studies of the AHE and SHE.
@@
This paper will stimulate many future studies of transport phenomena in periodically driven systems. First, our results allow the experimental detection of the spin current in periodically driven systems via the inverse SHE. The impact of our paper is not restricted to the SHE, but it will encourage future studies of other spintronics phenomena such as the spin Seebeck effect Saitoh-SSE ; Saitoh-SSE2 ; Bauer ; Adachi in periodically driven systems. This is because the inverse SHE is often used to convert the spin current into an electrical signal. Then, our results will provide a useful guideline when studying the Onsager reciprocal relations for the time-averaged charge and spin off-diagonal dc conductivities in the systems driven by CPL, LPL, or BCPL. Moreover, our general arguments and theory can be extended to the other transport coefficients including the ac conductivities in periodically driven systems. The extension to the transport coefficients in the non-linear regime is also an important future study.

### v2

@@ 6.1 Origin of the $\theta$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$
$\times[e^{-ik_{y}}(-1)^{n-m-(\beta-1)l}+e^{ik_{y}}],$ (101)
and $\mathcal{J}_{l}(u)$ is the $l$th-order Bessel function of the first kind. For simplicity, we restrict the Floquet indices $m$, $n$, and $l$ to be $1\leq m$, $n$, $l\leq 1$, which corresponds to the case at $n_{\textrm{max}}=1$. After some calculation, we obtain $I_{x}^{mn}(k_{x},u,\theta,\beta)$’s at $\beta=2$ and $3$,
$I_{x}^{mn}(k_{x},u,\theta,2)$ $\sim 2\delta_{n-m,0}\cos k_{x}\mathcal{J}_{0}(u)^{2}$
@@
$-2(e^{-i\theta}\delta_{n-m,2}+e^{i\theta}\delta_{n-m,-2})\sin k_{x}\mathcal{J}_{0}(u)\mathcal{J}_{1}(u)$
$+O(u^{3}),$ (102)
@@
$I_{x}^{mn}(k_{x},u,\theta,3)$ $\sim 2\delta_{n-m,0}\cos k_{x}\mathcal{J}_{0}(u)^{2}$
@@
$-2(e^{-i\theta}\delta_{n-m,2}+e^{i\theta}\delta_{n-m,-2})\cos k_{x}\mathcal{J}_{1}(u)^{2}$
$+O(u^{3}).$ (103)
@@
$I_{x}^{mn}(k_{x},u,\theta,2)$ $\sim\delta_{n-m,0}(2-u^{2})\cos k_{x}$
$-(\delta_{n-m,1}+\delta_{n-m,-1})u\sin k_{x}$
@@
$-(e^{-i\theta}\delta_{n-m,2}+e^{i\theta}\delta_{n-m,-2})u\sin k_{x}$
$+O(u^{3}),$ (104)
@@
$I_{x}^{mn}(k_{x},u,\theta,3)$ $\sim\delta_{n-m,0}(2-u^{2})\cos k_{x}$
$-(\delta_{n-m,1}+\delta_{n-m,-1})u\sin k_{x}$
@@
$-(e^{-i\theta}\delta_{n-m,2}+e^{i\theta}\delta_{n-m,-2})\frac{u^{2}}{2}\cos k_{x}$
$+O(u^{3}).$ (105)
@@
These equations can explain the $\theta$-independent $\sigma_{yx}^{\textrm{S}}$ for small $u$ because the terms for nm=0n-m=0 are primary (if they give the finite contribution). Although the terms for $n-m\neq 0$ are necessary for $\sigma_{yx}^{\textrm{C}}\neq 0$, these equations can also explain the $\theta$-independent $\sigma_{yx}^{\textrm{C}}$ for small $u$, at which the terms for $n-m=\pm 1$ are dominant. This is because in the small-$u$ region, the $\theta$-independent $u^{1}$ terms are more important than the $\theta$-dependent $u^{2}$ terms. For moderately large $u$, the $\theta$-dependent $u^{2}$ terms for $n-m=\pm 1$ may become non-negligible. Therefore, the $\theta$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ could be interpreted in terms of the $u$ and $\theta$ dependences of the light-induced corrections to the kinetic energy.
6.2 Comparisons with other relevant studies
@@
We now compare our results about the Onsager reciprocal relations with other relevant studies. As described in Sect. 1, there are several studies [24, 25, 26, 27] about the Onsager reciprocal relations in periodically driven systems. In two of them [24, 25], the time-periodic field was treated as perturbation. Meanwhile, in our theory, the time-periodic pump field has been treated nonperturbatively. In general, the time-periodic field should be treated nonperturbatively to describe periodically driven systems because its nonperturbative effects often change the electronic states drastically. Therefore, the systems studied in these papers [24, 25] are insufficient to discuss the standard periodically driven systems. Then, in another [26], the Onsager reciprocal relations about the matter or energy transport between a periodically driven system and one of the three reservoirs were studied. This is distinct from the transport phenomena within periodically driven systems. Therefore, except our previous paper [27], there has been no study of the Onsager reciprocal relations for any transport within periodically driven systems in which the nonperturbative effects of the driving field are taken into account. In that paper [27], we numerically studied the time-averaged charge off-diagonal dc conductivities $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{xy}^{\textrm{C}}$ in graphene driven by CPL or LPL and found that $\sigma_{yx}^{\textrm{C}}=-\sigma_{xy}^{\textrm{C}}$ or $\sigma_{yx}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}$ holds with CPL or LPL, respectively. However, we did not discuss the generality of these relations, i.e., we just showed that they are satisfied in our numerical results. Meanwhile, in this paper, we have made general arguments about the Onsager reciprocal relations for spin and charge transport within the periodically driven systems. These arguments are applicable to many periodically driven systems and extendable to others. Therefore, this work is the first systematic study of the Onsager reciprocal relations in periodically driven systems. Most importantly, this paper is the first work demonstrating the Onsager reciprocal relations for the spin transport in periodically driven systems.
@@
Next, we compare our results in the case of BCPL at $\beta=3$ with our previous study [30] for graphene driven by BCPL at $\beta=3$. (Although we have also studied the case at $\beta=2$ there [30], its results are not directly comparable to the results obtained in this paper due to the vital role of the valley degree of freedom in that case studied in Ref. [30].) There are three similarities between our results shown in Sect. 5.2.2 and those obtained in this previous study [30]: $\sigma_{yx}^{\textrm{C}}$ is almost independent of $\theta$ for small $u$ and depends on it for moderately large $u$; $\sigma_{yx}^{\textrm{C}}$ for small $u$ is dominated by the antisymmetric part; and the antisymmetric and symmetric parts are mixed for moderately large $u$ at $\theta=\frac{\pi}{2}$. Therefore, these properties may be characteristic properties of $\sigma_{yx}^{\textrm{C}}$ in systems driven by BCPL. Then, the difference is that the antisymmetric part of $\sigma_{yx}^{\textrm{C}}$ becomes almost vanishing for moderately large $u$ in Sr2RuO4 driven by BCPL at $\theta=0$ and $\pi$, whereas the antisymmetric part remains dominant even for moderately large $u$ in graphene driven by BCPL at these $\theta$’s [30].
@@
We turn to the comparisons with some theoretical studies [66, 67, 68] about transport properties with BCPL at $\beta=2$. In these studies, the effects of the BCPL on charge or spin transport have been analyzed in perturbation theories. First, our results about the effects of the SOC are similar to those on the charge and spin conductivities of another three-orbital electron system having the $z$ component of the SOC in the second-order perturbation theory against the BCPL field [67]. This similarity may be reasonable because the SOC is treated nonperturbatively in both cases. Then, our $\theta$-dependent $\sigma_{yx}^{\textrm{C}}$ is similar to the $\theta$-dependent charge conductivity obtained in the third-order perturbation theory against the BCPL field [68]. However, a critical value of $u$ above which $\sigma_{yx}^{\textrm{C}}$ depends on $\theta$ exists only in our case [see Fig. 12(a)]. This critical value may arise from the nonlinear $u$ terms of the light-induced corrections, as we have discussed in Sect. 6.1. Therefore, the crossover between the $\theta$-independent and $\theta$-dependent $\sigma_{yx}^{\textrm{C}}$ may be characteristic of nonperturbative effects of BCPL. This interpretation remains valid in graphene driven by BCPL [30]. Although the authors of Ref. [68] have claimed that a nonperturbative effect is also discussed, their discussions are insufficient because they have used an approximation for the Hamiltonian, $H_{0}[k+A(t)]\approx H_{0}(k)+A(t)\frac{\partial H_{0}(k)}{\partial k}$ [see Eq. (29) of Supplemental Material of Ref. [68]], which is valid only if $A(t)$, the vector potential of the BCPL, can be treated perturbatively. Namely, their theory can analyze only the perturbative effects of the BCPL. We do not use such an approximation; instead, we have analyzed the nonperturbative effects of BCPL in the standard Floquet linear-response theory [38, 8, 49, 11].
6.3 Experimental realization
@@
Finally, we comment on experimental realization of our results. We have supposed that our periodically driven open system can reach a nonequilibrium steady state due to the damping $\Gamma$. The AHE predicted in such a periodically driven open system [7] was experimentally observed [10]. Therefore, our time-averaged charge off-diagonal dc conductivities could be experimentally observed in the pump-probe measurements for periodically driven Sr2RuO4. In addition, our SHE could be detected via the inverse SHE [22, 23] because we have demonstrated that $\sigma_{yx}^{\textrm{S}}$ satisfies the Onsager reciprocal relation, which is similar to that in nondriven systems, and is dominated by the antisymmetric part. Note that $\tilde{\sigma}_{xy}^{\textrm{S}}$ could be experimentally observed by measuring the charge current perpendicular to the probe spin field [see Fig. 4(f)]. As the probe spin field, we can use, for example, the gradient of the Zeeman field, which could be experimentally realized [69]. In the case of Sr2RuO4, $u=0.1$ for $\Omega=6$ eV corresponds to $E_{0}\approx 15.4$ MV/cm, where we have used $u=ea_{\textrm{NN}}A_{0}=ea_{\textrm{NN}}E_{0}/\Omega$ and $a_{\textrm{NN}}\approx 0.39$ nm [70]. From an experimental point of view, the pump field of the order of $10$ MV/cm can be realized [71]. Therefore, it may be possible to observe the $\theta$-dependent $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ for moderately large $u$ as well as the $\theta$-independent $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ for small $u$. Although there is a possibility that the similar $\theta$-dependent $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{S}}$ will be realized even for smaller $\Omega$ and $E_{0}$, it is difficult to check this possibility due to the huge cost of the numerical calculations. Note that in the case of CPL [11], the $u$ dependence of the $\sigma_{yx}^{\textrm{S}}$ or $\sigma_{yx}^{\textrm{C}}$ obtained for $\Omega=6$ eV is similar to that obtained for smaller $\Omega$.
7 Conclusions
@@
Our arguments and numerical calculations have demonstrated that $\sigma_{yx}^{\textrm{S}}$ in the periodically driven systems satisfies the Onsager reciprocal relation. This means that the finite $\sigma_{yx}^{\textrm{S}}$ in the periodically driven systems can be indirectly observed by measuring the time-averaged dc conductivity of the inverse SHE. Therefore, our results open the way for detecting the spin current in periodically driven systems via the inverse SHE. Since such an indirect detection method has been widely used in many spintronics phenomena of nondriven systems [72], this achievement is useful to develop and observe many spintronics phenomena in periodically driven systems. Therefore, our results provide a vital step towards comprehensive understanding and further development of spintronics in periodically driven systems.
@@
Moreover, our arguments have resolved the contradictory statements about the Onsager reciprocal relation for the spin transport in nondriven systems [35, 36, 34, 37]. There is a previous study claiming the violation of the Onsager reciprocal relation for the spin transport [36], whereas there are other studies claiming its existence [35, 34, 37]. Furthermore, among the latter studies, the expression of the Onsager reciprocal relation is different: in one of them [35] the spin off-diagonal dc conductivity is dominated by the symmetric part, whereas in the others [34, 37] it is dominated by the antisymmetric part. As we have shown in Sect. 2.2.1, our results are consistent with the results of the last two studies [34, 37], i.e., the spin off-diagonal dc conductivity satisfies the Onsager reciprocal relation even with time-reversal symmetry and is dominated by the antisymmetric part. Therefore, our arguments support the validity of the interpretation of the inverse SHE [22, 23] as the existence of the spin current in nondriven systems.
@@
Then, our numerical calculations in the cases with BCPL have indicated that $\sigma_{yx}^{\textrm{C}}$ cannot necessarily be regarded as the anomalous Hall conductivity even if time-reversal symmetry is broken. In general, we can regard $\sigma_{yx}^{\textrm{C}}$ as the anomalous Hall conductivity if and only if $\sigma_{yx}^{\textrm{C}}$ is dominated by the antisymmetric part. In addition, time-reversal symmetry can be broken by BCPL [30]. As shown in Sect. 5.2.2, $\sigma_{yx}^{\textrm{C}}$ with BCPL for moderately large $u$ is almost vanishing or dominated by the symmetric part, although that for small $u$ is dominated by the antisymmetric part. This unusual property does not contradict the Onsager reciprocal relation because this relation in the case with BCPL restricts $\sigma_{yx}^{\textrm{C}}$ to neither the antisymmetric nor symmetric part generally (see Sect. 2.1.2). It may be due to the lack of a simple relation between the pump field of BCPL and its time-reversal counterpart. In contrast, the pump field of CPL has the simple relation $\mathbf{A}_{\textrm{LCPL}}(-t)=\mathbf{A}_{\textrm{RCPL}}(t)$; as a result, $\sigma_{yx}^{\textrm{C}}$ with CPL is restricted to the antisymmetric part (see Sect. 2.1.2). These results suggest that even if time-reversal symmetry can be broken by the pump field, it is highly required to check whether or not the main term of $\sigma_{yx}^{\textrm{C}}$ is given by the antisymmetric part in discussing the AHE. It is also necessary to check the main term of $\sigma_{yx}^{\textrm{S}}$ to discuss the SHE because in some cases such as the cases with BCPL its main term is not restricted by the Onsager reciprocal relation to either the antisymmetric or symmetric part. These suggestions are useful for future studies of the AHE and SHE.
@@
This paper will stimulate many future studies of transport phenomena in periodically driven systems. First, our results allow the experimental detection of the spin current in periodically driven systems via the inverse SHE. The impact of our paper is not restricted to the SHE, but it will encourage future studies of other spintronics phenomena such as the spin Seebeck effect [73, 74, 72, 75] in periodically driven systems. This is because the inverse SHE is often used to convert the spin current into an electrical signal. Then, our results will provide a useful guideline when studying the Onsager reciprocal relations for the time-averaged charge and spin off-diagonal dc conductivities in the systems driven by CPL, LPL, or BCPL. Moreover, our general arguments and theory can be extended to the other transport coefficients including the ac conductivities in periodically driven systems. The extension to the transport coefficients in the non-linear regime is also an important future study.